Undergraduate Software Engineering at RIT

Michael J. Lutz

Department of Computer Science, RIT

Abstract

This talk will focus on the progress to date in developing an undergraduate software engineering program at RIT. In the course of the presentation, the following issues will be addressed:

• What is software engineering?

• How does software engineering differ from traditional engineering?

• How does software engineering differ from computer science?

• What should be the goals of an undergraduate program in software engineering?

• What is the proposed curricular framework?

• What are the pragmatics and politics that have influenced the program’s development?

• What is planned for the immediate future?

Refreshments and cookies will be provided